Solve by using the quadratic formula - (x-3)(x+2)=5

Answer by checkley71(8403)

-(X-3)(X+2)=5
-(X^2-3X+2X-6)=5
-(X^2-X-6)-5=0
-X^2+X+1=0
X^2-X-1=0
X=(1+-SQRT[1^2-4*1*-1])/2*1
X=(1+-SQRT[1+4])/2
X=(1+-SQRT5)/2
X=(1+-2.236)/2
X=(1+2.236)/2
X=3.236/2
X=1.618 ANSWER.
X=(1-2.236)/2
X=-1.236)/2
X=-.618 ANSWER.